OpenStudy (123456789):

the included angle of the two sides of constant equal length s of an isosceles triangle is (theta) a) show that the area of the triangle is given by A = .5 s(squared) sin (theta). (Done) b) if theta is increasing at the rate of .5 radian per minute, find the rates of change of the area when (theta)= pi/6 and (theta)/3 c) explain why the rate of change of the area of the triangle is not constant even though d(theta)/dt is constant.

OpenStudy (anonymous):

\[ \frac{d \mathcal{A}}{d \theta} = s * \cos{\theta} * \frac{d \theta}{dt} \] and \frac{d \theta}{dt} is known, so it would be easy to find the answer for question "b", and for question "c", i think it would be better if you find it yourself.

OpenStudy (anonymous):

Sorry, I mean \[\frac{d\theta}{dt}\] is known

OpenStudy (anonymous):

Great explanation